Please help!! I need to find the equation of a line going through (6,-3) and perpendicular to the line y=3x + 4. Can you help me with this?
Answer by queenofit(120)

Remember, in perpendicular lines, the slope is the neg, reciprocal of the slope of the first line. In this case the slope of the first line is 3 so its negative reciprocal is -1/3. use this number as your slope for the perpendicular line.
now put that in the following formula
y-y1=m(x-x1)
If you are wondering about the 1's in the equation they indicate the first set of points you have. If you had 2 sets of points, then you would know which set to use.
for your problem put the points in and the slope (m) so...
y-(-3)=-1/3(x-6) following me?
y+3=-1/3x+2
y=-1/3x+2-3
y=-1/3x-1
do you see how I got the answer?